8) (ex,a -ax)Dx +&* YD, DO Y(0)=2

differential equations initial value
Answer #1

2-11 soli - Given differential Eqn ( +- 2x) + y 2 = 0 = 0 (%) => Tequ 0 is of the form 1. M dx + N dy =o where m= e 2x g²_2a- 2 = c ** - .. y eza_2x² = 20 gre za = 22 +222 e2u y = /2 cena + 2 2² e 2n - ② But 460] =2 7(0) = 2 in Eqn 09 in Egn (2) lie